What is the peak wavelength of light coming from a star with a temperature of 7,750 K? (Submit your answer in nanometers. Remember 1nm = 10-9 m)
Answer #1

Using the Wein's displacement law we have

(max)*T = 2898*10^-6

(max)*7750 = 2898*10^-6

(max) = 374*10^-9 m

so the peak wavelength is 374 nm